ok im struggleing with this so is every one else here is the question
a/0.6 = 3a + 2

Answer by nerdybill(7384)

a/0.6 = 3a + 2
multiply both sides by 0.6:
a = 0.6(3a + 2)
a = 1.8a + 1.2
-0.8a = 1.2
a = -1.5
.
To check, plug the answer back into the original equation:
a/0.6 = 3a + 2
-1.5/0.6 = 3(-1.5) + 2
-2.5 = -4.5 + 2
-2.5 = -2.5 (checks)